OpenStudy (anonymous):

d(x^2+1)--------dx(x^2-1)please solve this

OpenStudy (anonymous):

do know about quotient rule?

OpenStudy (anonymous):

\[\frac{ d }{ dx }(\frac{ x }{ y }) = \frac{ y*d/dx(x) - x*d/dx(y) }{ y^{2}}\]

OpenStudy (anonymous):

please buddy solve this if i know that why i ask you

OpenStudy (anonymous):

?

OpenStudy (anonymous):

--------what's this

OpenStudy (anonymous):

What is your question?